The adobe package^{∗}

Sebastian Rahtz S.Rahtz@elsevier.co.uk

1995/08/16

1 Introduction

This file contains suitable package files to load certain non-standard Adobe PostScript fonts. The font definition files and font metrics are available in both T1 and OT1 form in the companion CTAN fonts/psfonts collection.

2 The docstrip modules

1 ⟨∗garamond⟩

 2 \NeedsTeXFormat{LaTeX2e}
 3 \ProvidesPackage{garamond}[\filedate\space\fileversion\space
 4 Garamond PSNFSS2e package]
 5 \renewcommand{\rmdefault}{pgm}
 6 \renewcommand{\sfdefault}{pop}
 7 \renewcommand{\ttdefault}{pcr}

8 ⟨/garamond⟩ 9 ⟨∗basker⟩

10 \NeedsTeXFormat{LaTeX2e}
11 \ProvidesPackage{basker}[\filedate\space\fileversion\space
12 Baskerville PSNFSS2e package]
13 \renewcommand{\rmdefault}{pnb}

14 ⟨/basker⟩ 15 ⟨∗mtimes⟩

16 \NeedsTeXFormat{LaTeX2e}
17 \ProvidesPackage{mtimes}[\filedate\space\fileversion\space
18 Monotype Times PSNFSS2e package]
19 \renewcommand{\rmdefault}{mnt}

20 ⟨/mtimes⟩ 21 ⟨∗bembo⟩

22 \NeedsTeXFormat{LaTeX2e}
23 \ProvidesPackage{bembo}[\filedate\space\fileversion\space
24 Bembo PSNFSS2e package]
25 \renewcommand{\rmdefault}{pbb}
26 \renewcommand{\sfdefault}{pop}
27 \renewcommand{\ttdefault}{pcr}

28 ⟨/bembo⟩

∗This file has version number 1.0, last revised 1995/08/16.